# Renormalization group improved cosmology in the presence of a stiff matter era

### v1

@@ IV.3 Anti radiation and stiff matter era
We next consider that in the very early universe there was stiff-matter as well as anti-radiation fluid which indicates that the modified energy density takes the form
@@
with $\Omega_{\text{arad},*}$ denoting the energy-density corresponding to the anti-radiation part divided by the energy-density at the fixed time limit. It is still evident to implement the fact that $\Omega_{\text{s},*}>\Omega_{\text{arad},*}$ as the first term in the right hand side of the above equation shall dominate the second term for very small values of the scale factor. If we claim that $\varepsilon=0$, then eq.(59) gives that
@@
IV.4 Stiff matter, anti radiation, and dark matter
@@
V Flow improved cosmology after the fixed point regime
We shall now make use of eq.(s)(16,17) in the Friedmann equations to investigate the behaviour of the scale factor and energy density after the fixed point regime. Using the forms of $G(k)$, $\Lambda(k)$, and $\varepsilon$ from eq.(s)(16,17,20) and substituting it back in eq.(11), we can obtain the Friedmann equation after the fixed point regime upto $\mathcal{O}(k^{4})$ as
$\begin{split}H^{2}\simeq\frac{\Lambda_{0}}{3}~{}.\end{split}$ (71)
The above result leads to the solution
$a_{0}(t)\simeq e^{\sqrt{\frac{\Lambda_{0}}{3}}t}$ (72)
which indicate the accelerated expansion due to dark-energy in the current time. Now we shall start with few of the cases that we have discussed in section (IV).
V.1 Stiff matter, radiation, and dark matter era after fixed point regime
We continue with the model proposed in section (IV.2). In this regime, $a>a_{*}$ and as a result $\xi>1$. Using the form of $k$ from eq.(51) and substituting it back in eq.(70), we can write down the modified Friedmann equation outside of the fixed point regime as
We can further simplify the above relation by dropping the radiation contribution as well. Eq.(75) then can be simplified as
One can use the values of $g_{*}$ and $\lambda_{*}$ as $g_{*}\simeq 0.27$, and $\lambda_{*}\simeq 0.36$ [22]. Now as discussed in [22], $\nu$ and $\omega$ are both positive numbers. At the fixed point $k\simeq m_{pl}\propto\frac{1}{\sqrt{G_{0}}}$ (in natural units), which helps us to approximately estimate $\varepsilon_{*}$ from eq.(20) as
$\varepsilon_{*}\simeq\frac{\nu}{4\pi\omega G_{0}^{2}}.$ (78)
We can use the dark-matter parameter $\Omega_{\text{m},*}$ to be 0.1 as have we used in sections (IV.2,IV.4) to plot the scale factors and energy densities against time. Instead of using proper values for the parameter, one is quite safe with the consideration $\sigma_{\text{m}}<1$. For our consideration we have used $\sigma_{\text{m}}=0.1$. We now plot $\frac{a_{\text{m}}(t)}{a_{*}}$ against $H_{0}t$ for the current case against the case with only dark-energy contribution in Fig.(10).
Figure 10: Scale factor versus time plot after the fixed point regime.
Doing an integration of both sides of the above equation, we obtain
where $\sigma_{\text{r}}$ is defined as
We plot $\frac{a_{\text{r}}(t)}{a_{*}}$ against $\frac{a_{0}(t)}{a_{*}}$ versus $H_{0}t$ in Fig.(11) where we have set $\sigma_{\text{r}}=0.15$ and $H_{0}t_{*}\simeq 0.68$. Again we observe that a radiation dominated era results in a change in the long time behaviour of the accelerated expansion of the universe which is same as the case as has been observed for the matter dominated era. In the next subsection, we consider the behaviour of the scale factor after the fixed point regime when considering the case of section(IV.4) with an universe filled with stiff-matter+anti-radiation+dark-matter fluid. One important thing to remember about the scale factor plots is that before the fixed point regime one should not plot upto the fixed point which stands same for the current case. For the sake of continuity of the result, we have plotted upto and from the fixed point for section(IV) and section(V) respectively.
V.2 Dark matter, anti radiation, and stiff matter era after fixed point regime
In section (IV.4) we have investigated the early universe with an stiff matter+anti-radiation+dark matter era. We found out that the early time dynamics although is dominated mainly by the stiff matter era, the dynamics tends to deviate as an effect of the anti radiation era and the dark-matter era. Now, in this subsection we shall investigate the beyond fixed time implication for the same model (perturbative regime). With the identification of $k$ from eq.(67) one can write down the modified Friedmann equation from eq.(70) for the perturbative regime as (we get similar expression from eq.(73) where $*\Omega_{\text{rad,*}}$ is replaced by $*-\lvert\Omega_{\text{arad,*}}\rvert$)
$\begin{split}\frac{H^{2}}{H_{0}^{2}}\simeq 1+\left(\frac{2\nu}{\omega\Lambda_{0}}\right)\left(\frac{8\pi g_{*}\varepsilon_{*}}{\lambda_{*}}\right)^{\frac{1}{2}}\left[\frac{\Omega_{\text{m},*}}{(a/a_{*})^{3}}+\frac{\Omega_{\text{s},*}}{(a/a_{*})^{6}}-\frac{\lvert\Omega_{\text{arad,*}}\rvert}{(a/a_{*})^{4}}\right]^{\frac{1}{2}}-\frac{\nu G_{0}}{\Lambda_{0}}\left(\frac{8\pi g_{*}\varepsilon_{*}}{\lambda_{*}}\right)\left[\frac{\Omega_{\text{m},*}}{(a/a_{*})^{3}}+\frac{\Omega_{\text{s},*}}{(a/a_{*})^{6}}-\frac{\lvert\Omega_{\text{arad,*}}\rvert}{(a/a_{*})^{4}}\right]\end{split}$ (83)
@@ V.2 Dark matter, anti radiation, and stiff matter era after fixed point regime
$H_{0}(t-t_{*})\simeq\int_{1}^{\xi}\frac{d\xi}{\xi}\left(1+\frac{\nu}{\omega\Lambda_{0}}\left(\frac{8\pi\text{g}_{*}\varepsilon_{*}\Omega_{\text{m},*}}{\lambda_{*}}\right)^{1/2}\xi^{-3/2}-\frac{\nu\left|\Omega_{\text{arad},*}\right|}{2\omega\Lambda_{0}}\left(\frac{8\pi\text{g}_{*}\varepsilon_{*}}{\lambda_{*}\Omega_{\text{m}*}}\right)^{1/2}\xi^{-5/2}\right)^{-1}~{}.$ (86)
@@
$\begin{split}a(t)&\simeq a_{*}e^{H_{0}(t-t_{*})+\frac{2\varsigma_{\text{m}}}{3}-\frac{2\varsigma_{\text{arm}}}{5}}\\ \implies a(t)&\simeq a_{*}e^{H_{0}(t-\tau)}\end{split}$ (89)
@@
with the identification of $\tau=t_{*}-\frac{2}{3\text{H}_{0}}\left(\varsigma_{\text{m}}-\frac{3\varsigma_{\text{arm}}}{5}\right)$. Hence, it is clear to understand that the very late time dynamics completely agrees with the observed accelerated expansion of our Universe. As the plot of the scale factor have the exponential behaviour, we have not plotted $\frac{a(t)}{a_{*}}$ against $H_{0}t_{*}$ for the current case.
VI Polytropic equation of state and inflation
As we have seen in the earlier analysis, an inflationary phase of the universe cannot exist before the fixed time regime as in order for the universe go through inflation, one needs to have a phase where $a(t)$ grows rapidly with time. As the fixed point is before the Planck-time regime, it is quite straight forward to claim that inflation can only happen in the later part of the evolution which is after Planck time. Now previously we were getting accelerated expansion in the beyond fixed time regime but only for the existence of the cosmological constant which implies the current accelerated expansion. In our current analysis, we set $\Lambda_{0}=0$ in eq.(70). Hence, neglecting the $\mathcal{O}(k^{4})$ term, it is quite plausible to write the Friedmann equation in the beyond fixed point regime as
$H^{2}\simeq\frac{2\nu k^{2}}{3\omega}~{}.$ (90)
Before moving forward to the identification of $k$, we start by considering an equation of state as [34, 38, 39, 40]
where $0\leq\alpha\leq 1$, $n>0$, and $\rho_{P}=\frac{1}{G_{0}^{2}}$ denoting the Planck density. Substituting the above equation of state in the continuity equation eq.(7), one can obtain the form of the energy density as
where $a^{(c)}$ denotes an integration constant. It is important to remember that at the Planck time $\varepsilon\simeq\rho_{P}$ indicating that at this point $a\ll a^{(c)}$. This indicates that not only $a_{*}\neq a^{(c)}$ but also $a_{*}\ll a^{(c)}$. Comparing the above equation with eq.(20), we obtain
The above result is quite interesting as above result is identical to the one obtained in [34]. This serves as a consistency check for the renormalization group improved cosmological model with the standard cosmological model discussed in [34]. Using the identification $t_{\text{Pl}}=\frac{1}{\sqrt{G_{0}\rho_{p}}}$, the general solution of eq.(94) for $\alpha=n=1$ reads
$\begin{split}&\sqrt{\frac{8\pi}{3}}\frac{t-t_{\text{Pl}}}{t_{\text{Pl}}}=\\ &\frac{2}{3(1+\alpha)}\left[\sqrt{1+\xi_{c}^{3(1+\alpha)}}+\ln\left[\frac{\xi_{c}^{\frac{3(1+\alpha)}{2}}}{1+\sqrt{1+\xi_{c}^{3(1+\alpha)}}}\right]\right]\Biggr{\rvert}_{\tilde{\xi}_{c}}^{\xi_{c}}\end{split}$ (95)
where $\xi_{c}$ and $\tilde{\xi}_{c}$ are defined as
$\xi_{c}=\frac{a}{a^{(c)}}~{},~{}~{}\tilde{\xi}_{c}=\frac{a_{*}}{a^{(c)}}~{}.$ (96)
As $a_{*}\ll a^{(c)}$ or $\tilde{\xi}_{c}\ll 1$ one can recast eq.(95) as
$\begin{split}&\frac{3(1+\alpha)}{2}\sqrt{\frac{8\pi}{3}}\frac{t-t_{*}}{t_{\text{Pl}}}\simeq\sqrt{1+\xi_{c}^{3(1+\alpha)}}\\ &-1+\ln\left[\left(\frac{a(t)}{a_{*}}\right)^{\frac{3(1+\alpha)}{2}}\right]-\ln\left[\frac{1+\sqrt{1+\xi_{c}^{3(1+\alpha)}}}{2}\right]\end{split}$ (97)
The fundamental difference of this result with the one presented in [34] is that in our case the solution is not valid before the Planck time which forces us to implement a lower bound of integration for the time part to be $t_{\text{Pl}}$ and the scale factor part to be $a_{*}$. For $a\gg a^{(c)}$, we arrive at the solution
$H^{2}=\frac{2\nu k^{2}}{3\omega}-\frac{\nu G_{0}k^{4}}{3}~{}.$ (100)
From eq.(s)(16,17), we can see that both $\omega$ and $\nu$ are correction parameters in the late time expansion of $G(k)$ and $\Lambda(k)$ respectively. Again $G_{0}^{2}$ itself is a very small quantity and as a result we can treat $\frac{\pi\rho_{p}\omega^{2}G_{0}^{2}}{\nu}$ as a small quantity and define it by $\eta$. Hence, we can recast eq.(101) (with the assumption that $\eta\ll 1$) as
$\begin{split}\sqrt{\frac{8\pi}{3}}\frac{t-t_{\text{Pl}}}{t_{\text{Pl}}}=&\frac{1}{3}\left(\sqrt{1+\xi_{c}^{6}}-\sqrt{1+\tilde{\xi}_{c}^{6}}\right)+(1-\eta)\ln\left[\nicefrac{{\xi_{c}}}{{\tilde{\xi}_{c}}}\right]\\ &-\frac{1-\eta}{3}\ln\left[\frac{\sqrt{1+\xi_{c}^{6}}+1}{\sqrt{1+\tilde{\xi}_{c}^{6}}+1}\right]~{}.\end{split}$ (104)
Here, $\eta=\frac{\pi\rho_{P}\omega^{2}G_{0}^{2}}{\nu}$ is a purely quantum gravitational correction and we find that the quantum gravitational correction amplifies the inflation. Instead of substituting the values of the individual parameters we consider an overall small value for $\eta$ and plot it against time where we compare the general case in [34] along with the quantum gravity modified inflationary scenario in Fig.(12).
Figure 12: Plot of the dimensionless scale factor versus the dimensionless time for $\eta=0$ against the case with $\eta=0.05$.
From Fig.(12), it is straightforward to check that the quantum gravity correction induces a higher rate of inflation corresponding to the case when $\eta=0$. The important thing to observe is that the value of $\eta$ is quite small for such a significant amount of change in the rate of inflation. This also helps the current models of inflation as the time of inflation reduces compared to the classical analogue. Now we shall investigate the general structure of the energy density required for having an inflationary model which can be thought of as a back calculation. We assume that the energy density has the form given by
where $\kappa,~{}j,~{}l\in\mathbb{R}$. The above form of the energy density can be motivated by directly taking an identification of the momentum scale $k$ in terms of the scale factor $a(t)$ as
$k=\left(\frac{4\pi\omega G_{0}\varepsilon_{P}}{\nu}\right)^{\frac{1}{2}}\left(1+\kappa\left[\frac{a}{a^{(c)}}\right]^{j}\right)^{-\frac{l}{2}}~{}.$ (107)
Using the form of $k$ in eq.(20), we get back the energy density in eq.(106). Substituting the form of the energy density from eq.(106 in the continuity equation in eq.(7), we can obtain the equation of state as
As in [34] as well as in our analysis the inflationary behaviour is dominant in the early phase when $a\ll a^{c}$, then only $\varepsilon\sim\varepsilon_{P}$ which is a constant. This leads to $k\propto\sqrt{\varepsilon_{P}}$ and $H=\text{constant}$. As a result, one obtains an inflationary behaviour when $a\ll a^{(c)}$. One can also propose an energy density of the form $\varepsilon=\varepsilon_{P}\left(1\pm\kappa^{\prime}\left(\frac{a}{a^{(c)}}\right)^{j^{\prime}}\right)^{l^{\prime}}$ [^6] For the mentioned form of the energy density the equation of state simply reads $P=-\left(1+\frac{j^{\prime}l^{\prime}}{3}\right)\varepsilon+\frac{j^{\prime}l^{\prime}\varepsilon}{3}\left(\frac{\varepsilon}{\varepsilon_{P}}\right)^{-\frac{1}{l^{\prime}}}$. If $j^{\prime}<0$ and $l^{\prime}>0$ such that $j^{\prime}l^{\prime}<-3$ then we get non-trivial equation of states where the energy density is of the form $\varepsilon=\varepsilon_{P}\left(1\pm\frac{\kappa^{\prime}}{\left(\frac{a}{a^{(c)}}\right)^{|j^{\prime}|}}\right)^{l^{\prime}}$. Now if one takes $j^{\prime}>0$ and $l^{\prime}<0$ such that $j^{\prime}l^{\prime}<-3$ then we get back the form similar to eq.(106). kind of structure with the only requirement being $\varepsilon\sim\text{constant}$ when $a\ll a^{(c)}$.
VII Conclusion
In this paper, we consider a cosmological model where the universe has gone through a stiff matter era. Instead of using the standard cosmological model, we make use of the renormalization group improved cosmology by considering the flow of the Newton’s gravitational constant as well as the cosmological constant with the momentum scale of the universe $k$. The $k$ dependence of $G(k)$ and $\Lambda(k)$ are different when one considers the epoch the universe before the fixed point regime and when one considers after the fixed point regime. At first we investigate the existence of a stiff-matter fluid before the fixed point regime along with several combinations of a dark matter, radiation, and anti-radiation era. To do so we identify the momentum scale of the universe $k$ with the scale factor via comparing the $k$ dependent form of the energy densoty before fixed point regime and scale factor dependent form of the energy density. We have then substituted the $a(t)$ dependent form of $k$ in the Friedmann equation and obtained a time-dependent solution of the scale factor. Plotting the dimensionless scale factor with change in the dimensionless time, we find out that for $a(t)\ll a_{*}$ the stiff matter era dominates and results in a general accelerated expansion of the universe upto the fixed point regime. Although we find out in section (IV.4), that a stiff matter era along with just anti-radiation will lead to several unexpected behaviour which we argued will be the same for a stiff matter together with anti-dark matter component as well. In order get a good behaviour of the evolution of the universe before the fixed point regime, one needs to consider a stiff matter+anti-radiation+dark matter combination or a stiff matter+radiation+anti-dark matter combination. One important aspect to note that unlike [34], the current model will break down for a anti-stiff matter component consideration as the starting energy density then will be zero which is in contradiction to the renormalization group improved approach at very high energy scales. We also argued that before fixed time, it is impossible to have an inflation era in the renormalization group framework. Next we have investigated the beyond fixed point regime for a stiff matter+radiation+dark matter combination, stiff matter+radiation combination, and stiff matter+anti-radiation+dark matter combination. We observe that in this regime where $a>a_{*}$, the dark matter part has the subleading order contribution where the leading order contribution to the accelerated expansion of the universe comes from the current value of the cosmological constant. In the absence of a dark matter component the radiation part has the next sub-leading contribution towards the evolution of the scale factor. Finally, in section(VI), we have used the polytropic equation of state introduced in [34, 38, 39, 40] and observe that in the beyond fixed point regime one can get inflation and inflation dynamics is similar to the one observed in [34]. The important observation is that the inflationary behaviour comes out from the $k^{2}$ order correction term of the Friedmann equation in eq.(70). This is an important result in our work. We have also computed quantum gravity corrections to the inflationary behaviour of the scale factor which indicates that inflation is enhanced by such effects. Finally, we have proposed the general class of equations of state that results in an inflationary model in the renormalization group improved cosmology as well as the general cosmological model discussed in [34].

### v3

@@ A.1 Anti radiation and stiff matter era before the Planck time
Here, we discuss that in the very early universe there was stiff-matter era followed by an anti-radiation fluid which indicates that the modified energy density takes the form
@@
with $\Omega_{\text{arad},*}$ denoting the energy-density corresponding to the anti-radiation part divided by the energy-density at the fixed time limit. The primarly problem of working with negative energy densities is that it may lead to the viloation of local energy conditions. As has been discussed in deltails in several extensions of quantum field theories QFTViolation ; QFTViolation2 ; QFTViolation3 ; QFTViolation4 ; QFTViolation5 ; QFTViolation6 ; QFTViolation7 , relativistic energy conditions only need to satisfy globally. This allows one for small local violations as has also been discussed in Patla . It is therefore possible that before the Planck time there had been a component with negative energy density which had contributed to the expansion of the universe for a very short period of time. It is still important from an analytical perspective to implement beforehand the fact that $\Omega_{\text{s},*}>\Omega_{\text{arad},*}$ as the first term in the right hand side of the above equation shall dominate the second term for very small values of the scale factor. If we claim that $\varepsilon=0$, then eq.(105) gives that
@@
A.2 Stiff matter, anti radiation, and dark matter
@@
A.3 Dark matter, anti radiation, and stiff matter era beyond the Planck time
In section (A.2) we have investigated the early universe with an stiff matter+anti-radiation+dark matter era. We found out that the early time dynamics although is dominated mainly by the stiff matter era, the dynamics tends to deviate as an effect of the anti radiation era and the dark-matter era. Now, in this subsection we shall investigate the beyond fixed time implication for the same model (perturbative regime). With the identification of $k$ from eq.(113) one can write down the modified Friedmann equation from eq.(54) for the perturbative regime as (we get similar expression from eq.(57) where $*\Omega_{\text{rad,*}}$ is replaced by $*-\lvert\Omega_{\text{arad,*}}\rvert$)
$\begin{split}\frac{H^{2}}{H_{0}^{2}}\simeq&1+\left(\frac{2\nu}{\omega\Lambda_{0}}\right)\left(\frac{8\pi g_{*}\varepsilon_{*}}{\lambda_{*}}\right)^{\frac{1}{2}}\left[\frac{\Omega_{\text{m},*}}{(a/a_{*})^{3}}+\frac{\Omega_{\text{s},*}}{(a/a_{*})^{6}}-\frac{\lvert\Omega_{\text{arad,*}}\rvert}{(a/a_{*})^{4}}\right]^{\frac{1}{2}}-\frac{\nu G_{0}}{\Lambda_{0}}\left(\frac{8\pi g_{*}\varepsilon_{*}}{\lambda_{*}}\right)\left[\frac{\Omega_{\text{m},*}}{(a/a_{*})^{3}}+\frac{\Omega_{\text{s},*}}{(a/a_{*})^{6}}-\frac{\lvert\Omega_{\text{arad,*}}\rvert}{(a/a_{*})^{4}}\right]\end{split}$ (116)
@@ A.3 Dark matter, anti radiation, and stiff matter era beyond the Planck time
$H_{0}(t-t_{*})\simeq\int_{1}^{\xi}\frac{d\xi}{\xi}\left(1+\frac{\nu}{\omega\Lambda_{0}}\left(\frac{8\pi\text{g}_{*}\varepsilon_{*}\Omega_{\text{m},*}}{\lambda_{*}}\right)^{1/2}\xi^{-3/2}-\frac{\nu\left|\Omega_{\text{arad},*}\right|}{2\omega\Lambda_{0}}\left(\frac{8\pi\text{g}_{*}\varepsilon_{*}}{\lambda_{*}\Omega_{\text{m}*}}\right)^{1/2}\xi^{-5/2}\right)^{-1}.$ (119)
@@
$\begin{split}a(t)&\simeq a_{*}e^{H_{0}(t-t_{*})+\frac{2\varsigma_{\text{m}}}{3}-\frac{2\varsigma_{\text{arm}}}{5}}\end{split}$ (122)
@@
where we can express the above equation as $a(t)\simeq a_{*}e^{H_{0}(t-\tau)}$ when $\tau\equiv t_{*}-\frac{2}{3\text{H}_{0}}\left(\varsigma_{\text{m}}-\frac{3\varsigma_{\text{arm}}}{5}\right)$. Hence, it is clear to understand that the very late time dynamics completely agrees with the observed accelerated expansion of our Universe. As the plot of the scale factor have the exponential behaviour, we have not plotted $\frac{a(t)}{a_{*}}$ against $H_{0}t_{*}$ for the current case.
Acknowledgement
G. G. extends his heartfelt gratitude to CSIR, Govt. of India for funding this research.